OpenStudy (anonymous):

Model the function rule y = -2x - 3 with a table of values.

OpenStudy (mathstudent55):

 For each value of x, insert it in the equation, and evaluate y. For example, for x = -3, you do: y = -2(-3) - 3 = 6 - 3 = 3